NEED ASAP IN 10 MINS PLEASE SHOW CALCULATIONS!!!!!!Sunrise Corporation issued $900,000 of 6%, 10-year bonds on January 1, 2021, for $777,690. This price provided a yield of 8%
Answer #1

Interest paid on June 30 = 900000*6%*6/12 = 27000

Interest expense on June 30 = 777690*8%*6/12 = 31108

Discount amortization = 31108-27000 = 4108

Interest paid on Oct 31 = 900000*6%*4/12 = 18000

Interest expense on Oct 31 = (777690+4108)*8%*4/12 = 20848

Total interest expense = 31108+20848 = 51956